An innovative restaurate owns and operates a dozen "Ultimate Low-Carb restaurants in Ontario. His signature item is a cheese-encrusted beef medallion wrapped in lettuce. Sales (X, in millions of dollars) is related to Profits (Y, in hundreds of thousands of dollars) by the regression equation Y 8.21 + 0.76 X. What is your forecast of profit for a store with sales of $40 million? $50 million?
Answer #1

Given, Y =8.21 + 0.76*X where Y=Profit in 100000 and X=Sales in 1000000

a) Sales of 40 million ->X=40

Y=8.21 + 0.76*40=38.61

Hence, Profits are $3861000

b) Sales of 50 million ->X=50

Y=8.21 + 0.76*50=38.61

Hence, Profits are $4621000.